A cookie recipe for 60 cookies calls for 4 cups of flour. How much flour is needed to make 90 cookies?


1
Expert's answer
2022-02-21T17:56:37-0500

Let x be the number of cookies we get for 1 cup of flour. Then,

"\begin{array}{|c|c|}
\hline
\text{Number of cup of flour}& \text{Number of cookies}\\
\hline
4&60\\
1&x\\
\hline
\end{array}\\"


"\therefore x = \dfrac{60}{4}=15" i.e., 1 cup of flour makes 15 cookies.

Let y be the number of cups of flour required to make 90 cookies.


"\begin{array}{|c|c|}
\hline
\text{Number of cup of flour}& \text{Number of cookies}\\
\hline
1&15\\
y&90\\
\hline
\end{array}\\"


"i.e., 15y = 90\Rightarrow y = \dfrac{90}{15}=6". Hence we require 6cups of flour to make 90 cookies.